What is the mass percent composition of manganese and oxygen in the oxide?
Given: 3.144g of manganese reacts with oxygen to produce 6.348g of manganese oxide.
1
Expert's answer
2018-07-25T08:02:26-0400
3,144 gr 6,348 gr

xMn + yO =˃ MnxOy

55x 55x +16y

3,144×(55x+16y)=6,348 × 55x

172,92 x + 50,304y = 349,14x

50,304y=176,22x

y=3,5x

x=2

y=7

Answer: Mn2O7